OpenStudy (anonymous):

The volume of a hexagonal prism is 6 ft3. What is the height of the prism if the base area is 2 ft3? 3ft 12 ft 18 ft 3 ft

OpenStudy (anonymous):

it didnt turn out the way it was supposed to . dont answer :C

jimthompson5910 (jim_thompson5910):

Volume = (Area of the base)*(Height) V = Ah 6 = 2h 6/2 = h 3 = h h = 3 So the height is 3 feet

OpenStudy (anonymous):

good Question. :)

OpenStudy (anonymous):

are you sure ? because my question didnt turn out like it was supposed to

jimthompson5910 (jim_thompson5910):

I'm very sure. If the volume is 6 cubic feet and the area of the base is 2 cubic feet, then the height is 3 feet.

OpenStudy (anonymous):

it didnt type right . when i pasted the answer the equation left :C

jimthompson5910 (jim_thompson5910):

That's alright. I understood it just fine. If you want, you can type "cubic feet" as "ft^3". The two are interchangeable.

hero (hero):

I think she's trying to say that she posted the question incorrectly

OpenStudy (anonymous):

The volume of a hexagonal prism is 6 ft\[\sqrt{?}\]3. What is the height of the prism if the base area is 2 ft\[\sqrt{?}\] 3

OpenStudy (anonymous):

there it is posted correctly

jimthompson5910 (jim_thompson5910):

Oh so the 3s are in the square roots?

OpenStudy (anonymous):

yes -.-

jimthompson5910 (jim_thompson5910):

If so, then... V = Ah 6*sqrt(3) = 2*sqrt(3)h (6*sqrt(3))/(2*sqrt(3)) = h 3 = h h = 3 and we get the same answer

OpenStudy (anonymous):

thanks !! :3

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 3 centimeters and a base that is an equilateral triangle with sides of 4 centimeters? 24 cm3 cm3 cm3 48 cm3 can you help me with this one too ?

OpenStudy (anonymous):

the 2 middle answers are 8squared3 and 12squared3

jimthompson5910 (jim_thompson5910):

With an equilateral triangle, if the base is x units, then half the base is x/2 units. The height is then (x/2)*sqrt(3) Therefore, the area is then A = (bh)/2 A = ((x/2)*(x/2)*sqrt(3))/2 A = (x^2/4)*sqrt(3) In this case, x = 4, so the area of the equilateral triangle is A = (4^2/4)*sqrt(3) A = 4*sqrt(3) Now let's find the volume Volume = (Area of base)*(Height) V = Ah V = 4*sqrt(3)*3 V = 12*sqrt(3) So the answer is choice C

OpenStudy (anonymous):

thanks ! your the best C:

jimthompson5910 (jim_thompson5910):

You're welcome

OpenStudy (anonymous):

yayy i got an 80% xD